OpenStudy (alois_trancy):

Help Again Please XD Will Fan And Medal! Given f(x) = 3x – 5, what is f(–4)? A. -6 B. 7 C. 1/3 D. -17

OpenStudy (baru):

3x - 5 replace the letter x with (-4) can you do that?

OpenStudy (alois_trancy):

Yeah Thanks :3

OpenStudy (baru):

what do you get?

OpenStudy (alois_trancy):

I'm Figuring It Out

OpenStudy (alois_trancy):

-17

OpenStudy (baru):

yep :) i hope you aren't guessing :P

OpenStudy (alois_trancy):

I Promise I'm not :3 Could you help me with a few others? If Not That's okay! :)

OpenStudy (baru):

i'll help :) but you have to work it out yourself

OpenStudy (alois_trancy):

Alrighty! One Second

OpenStudy (alois_trancy):

What is the slope of the graph of 8x – 4y = 11? A. 2 B. 1/2 C. -2 D. -1/2

OpenStudy (baru):

first you have to put it in this form y= mx + c can you try that?

OpenStudy (alois_trancy):

I think so...

OpenStudy (baru):

what do you get?

OpenStudy (alois_trancy):

-2? If not, I did the math wrong..

OpenStudy (baru):

what did you get when you put it in the form y=mx + c ?

OpenStudy (alois_trancy):

One sec...

OpenStudy (alois_trancy):

2

OpenStudy (alois_trancy):

@baru

OpenStudy (baru):

guessing aren't you :P sorry cant give you the answer, i can only tell you the steps to follow.

OpenStudy (alois_trancy):

I will let you know that no, i'm not guessing, you can think what you want to but the truth is the truth....